As a technician in a large pharmaceutical research firm, you need to produce 150.mL of 1.00 M potassium dihydrogen phosphate buffer solution of pH = 6.79. The pKa of H2PO4? is 7.21.

You have the following supplies: 2.00 L of 1.00 M KH2PO4 stock solution, 1.50 L of 1.00 M K2HPO4 stock solution, and a carboy of pure distilled H2O.

How much 1.00 M KH2PO4 will you need to make this solution?

Express your answer to three significant digits with the appropriate units.
Answer #1

We use Henderson-Hasselbalch equation.

pH= pKA + log [Base] / [Acid]

6.79 = 7.21 + log [HPO4] / [H2PO4]

log [HPO4] / [H2PO4] = -0.42

[HPO4] / [H2PO4] = 0.380 / 1.00

But, we know that [HPO4] + [H2PO4] = 1.00 M

[HPO4] = 1.00 - [H2PO4]

0.380 = [1.00 - [H2PO4] ] / [H2PO4]

0.380 [H2PO4] = 1.00 - [H2PO4]

1.380 [H2PO4] = 1.00

[H2PO4] = 0.725 M

We have 2.00 L of 1.00 M KH2PO4 stock solution. We need 150.0 mL of 0.725 M

Thus, M1V1 = M2V2

(1.00)(V1) = (0.725) (150)

V1 = 108.75 mL = 109 mL

Thus, we need 109 mL of 1.00 M KH2PO4 is needed.